Duffy Dog Mills is a division of Iowa Woolen Products. For the most recent year, Duffy Dog had net income of $19,400,000. Included in income was interest expense of $1,460,000. The operation’s tax rate is 40 percent. Total assets of Duffy Dog Mills are $237,000,000, current liabilities are $51,100,000, and $37,000,000 of the current liabilities are noninterest bearing.

Calculate NOPAT, invested capital, and ROI for Duffy Dog Mills. (Round ROI to 2 decimal places, e.g. 15.25%.)

NOPAT $
Invested capital $
ROI %
Answer #1

NOPAT = (Net Income/(1-Tax rate) + Interest Expense)(1-Tax rate)

= (19,400,000/0.6+ 1,460,000)(1-40%)

= $20,276,000

Invested Capital = Total Assets – Non interest bearing liabilities

= 237,000,000-37,000,000

= $200,000,000

ROI = NOPAT/Invested capital

= 20,276,000/200,000,000

= 10.138%

i.e. 10.14%